OpenStudy (anonymous):

Help with parameterization: so I am trying to model the second hand of a clock. What I know: the radius of the clock face is 5. With 60 minutes on the clock and 360 degrees on a circle, each position is 6 degrees apart. The xcoord is rcos(theta). The y coord is rsin(theta).

OpenStudy (dumbcow):

\[x(t) = 5\cos(\frac{\pi}{2} - \frac{\pi}{30} t)\] \[y(t) = 5\sin(\frac{\pi}{2} - \frac{\pi}{30} t)\] this models a second hand on a clock where t=0 is at 12 o clock